# Safety with Agency: Human-Centered Safety Filter with Application to AI-Assisted Motorsports

### v1

@@ Abstract
We propose a human-centered safety filter (HCSF) for shared autonomy that significantly enhances system safety without compromising human agency. Our HCSF is built on a neural safety value function, which we first learn scalably through black-box interactions and then use at deployment to enforce a novel quality control barrier function (Q-CBF) safety constraint. Since this Q-CBF safety filter does not require any knowledge of the system dynamics for both synthesis and runtime safety monitoring and intervention, our method applies readily to complex, black-box shared autonomy systems. Notably, our HCSF’s CBF-based interventions modify the human’s actions minimally and smoothly, avoiding the abrupt, last-moment corrections delivered by many conventional safety filters. We validate our approach in a comprehensive in-person user study using Assetto Corsa—a high-fidelity car racing simulator with black-box dynamics—to assess robustness in “driving on the edge” scenarios. We compare both trajectory data and drivers’ perceptions of our HCSF assistance against unassisted driving and a conventional safety filter. Experimental results show that 1) compared to having no assistance, our HCSF improves both safety and user satisfaction without compromising human agency or comfort, and 2) relative to a conventional safety filter, our proposed HCSF boosts human agency, comfort, and satisfaction while maintaining robustness.
@@ I Introduction
  • We introduce, to the best of our knowledge, the first fully model-free control barrier function (CBF) safety filter. We learn a neural safety value function through interactions with a black-box system and, at deployment, enforce a novel quality control barrier function (Q-CBF) safety constraint without any knowledge of system dynamics (e.g., control affine model). Both the synthesis and deployment of our Q-CBF are scalable to high-dimensional systems and do not require any knowledge of their dynamics.
@@ IV Smooth Human-Centered Safety Filter for Shared Autonomy
Conventional CBF safety filter methods similar to (9) typically require knowledge of the system’s dynamics [32, 10, 13, 15, 58, 31], even when using learned barrier functions [20, 33, 59, 60, 61, 62]. This requirement arises for one or both of the following reasons: 1) either a full-order or a simplified dynamical model of the system is utilized to synthesize CBF candidates, and 2) knowledge of the dynamics (e.g., control affine model) is leveraged at runtime to enforce the CBF safety constraint within an optimal control problem (OCP) (e.g., quadratic program). Some works have explicitly aimed to build and deploy model-free CBF safety filters, but they often fall short of being fully model-free—resorting to simplified, low-dimensional models of the system dynamics [63], requiring a velocity tracking controller alongside a simplified kinematic model [64, 65], or necessitating a predefined fallback policy (i.e., an evading maneuver) [66]. Moreover, while learning a CBF in latent space can be effective for partially observable systems [67], it still requires the control affinity assumption. Such reliance on knowing the system dynamics and the deployment environment can significantly limit the applicability of CBF s in complex, real-world scenarios where the dynamical model is often unknown and should be treated as a black-box.
@@ V Application of HCSF to High-Speed Racing
Our HCSF design in Section IV assumes the knowledge of the state–action safety value function ${\mathcal{Q}}(\cdot,\cdot)$. However, directly solving (6) for high-dimensional systems is intractable—a manifestation of the notorious “curse of dimensionality.” In this section, we leverage recent advances in safety RL to address this challenge and learn both the ${\mathcal{Q}}$-function and the best-effort fallback policy in a high-fidelity racing environment. Specifically, we begin by describing the environment setup and then detail the observation space, action space, and safety margin function. Next, we outline our multi-phase training pipeline, which includes warmup and initialization phases that expedite learning by frequently exposing the agent to “dangerous” states. Finally, we discuss how we integrate safety filters with visual cues in an effort to enhance the AI’s transparency.
@@ V-F Visual Cues
During our user study, all participants are shown a color-coded reference path which is a series of green and red arrowheads on the track (Fig. 2). Green arrowheads indicate acceleration, and red arrowheads advise deceleration. This visual aid allows drivers to navigate effectively even without expertise in racing.
@@ VI-A Baselines
For LRSF, we follow the definition in (7). This filter only intervenes when the system reaches the boundary of the safe set and does not consider the human operator’s input during interventions. We anticipate that such an abrupt approach which disregards human intent may confuse drivers and undermine comfort, thus increasing the risk of automation surprise. In contrast, our HCSF solves the optimization problem (11) to minimally modify the human action while still satisfying the Q-CBF constraint, which we hypothesize will enhance human agency, comfort, and overall satisfaction compared to LRSF. To ensure a fair comparison between our HCSF and LRSF, we employ the same neural approximation of the state–action safety value function ${\mathcal{Q}}(\cdot,\cdot)$ for both value-based safety filters. This ensures that both filters rely on the same safety monitor, while their interventions may differ (7), (11). In addition, each filter’s visual cues follow the same proportionality rule, maintaining consistency in how interventions are conveyed to human drivers.

### v2

@@ Abstract
We propose a human-centered safety filter (HCSF) for shared autonomy that significantly enhances system safety without compromising human agency. Our HCSF is built on a neural safety value function, which we first learn scalably through black-box interactions and then use at deployment to enforce a novel state–action control barrier function (Q-CBF) safety constraint. Since this Q-CBF safety filter does not require any knowledge of the system dynamics for both synthesis and runtime safety monitoring and intervention, our method applies readily to complex, black-box shared autonomy systems. Notably, our HCSF’s CBF-based interventions modify the human’s actions minimally and smoothly, avoiding the abrupt, last-moment corrections delivered by many conventional safety filters. We validate our approach in a comprehensive in-person user study using Assetto Corsa—a high-fidelity car racing simulator with black-box dynamics—to assess robustness in “driving on the edge” scenarios. We compare both trajectory data and drivers’ perceptions of our HCSF assistance against unassisted driving and a conventional safety filter. Experimental results show that 1) compared to having no assistance, our HCSF improves both safety and user satisfaction without compromising human agency or comfort, and 2) relative to a conventional safety filter, our proposed HCSF boosts human agency, comfort, and satisfaction while maintaining robustness.
@@ I Introduction
  • We introduce, to the best of our knowledge, the first fully model-free control barrier function (CBF) safety filter. We learn a neural safety value function through interactions with a black-box system and, at deployment, enforce a safety constraint based on a novel state–action control barrier function (Q-CBF) without any knowledge of system dynamics (e.g., control affine model). Both the synthesis and deployment of our Q-CBF are scalable to high-dimensional systems and do not require any knowledge of their dynamics.
@@ IV Smooth Human-Centered Safety Filter for Shared Autonomy
Conventional CBF safety filter methods similar to (9) typically require knowledge of the system’s dynamics [32, 10, 13, 15, 58, 31], even when using learned barrier functions [20, 33, 59, 60, 61, 62]. This requirement arises for one or both of the following reasons: 1) either a full-order or a simplified dynamical model of the system is utilized to synthesize CBF candidates, and 2) knowledge of the dynamics (e.g., control affine model) is leveraged at runtime to enforce the CBF safety constraint within an optimal control problem (OCP) (e.g., quadratic program).
While some works explicitly aimed to build and deploy model-free CBF safety filters, they have so far fallen short of being fully model-free—relying on some combination of simplified models of the system dynamics [63, 64, 65], predefined low-level controllers [64, 65], and handcrafted fallback policies (e.g., evading maneuvers) [66]. Additionally, recent efforts in learning a CBF for latent state representations have proven to be effective for partially observable systems, but they still require a control affine dynamical model [67]. Such reliance on knowledge of the system dynamics and the deployment environment can significantly limit the applicability of CBF s in complex, real-world scenarios where the dynamical model is often unknown and should be treated as a black-box. On the other hand, a model-free algorithm for learning a policy together with a barrier certificate was proposed recently [68], but it cannot be used to build a safety filter because the learned policy must be deployed at all times. Finally, we acknowledge a preprint reporting concurrent efforts toward a model-free state–action CBF safety filter [69]. However, it addresses a finite-horizon safety problem and enforces safety at runtime like a smooth least restrictive safety filter [70] rather than a CBF one, fundamentally differing from our work in both mathematical formulation and enforcement of safety.
@@ V Application of HCSF to High-Speed Racing
Our HCSF design in Section IV assumes the knowledge of the state–action safety value function ${\mathcal{Q}}(\cdot,\cdot)$. However, directly solving (6) for high-dimensional systems is intractable—a manifestation of the notorious “curse of dimensionality.” In this section, we leverage recent advances in safety RL to address this challenge and learn both the ${\mathcal{Q}}$-function and the best-effort fallback policy in a high-fidelity racing environment. Specifically, we begin by describing the environment setup and then detail the observation, action, safety margin function, and episode termination conditions. Next, we outline our multi-phase training pipeline, which includes warmup and initialization phases that expedite learning by frequently exposing the agent to “dangerous” states. Finally, we discuss how we integrate safety filters with visual cues in an effort to enhance the AI’s transparency.
@@ V-F Visual Cues
During our user study, all participants are shown a color-coded reference path which is a series of green and red arrowheads on the track (Fig. 2). Green arrowheads indicate acceleration, and red arrowheads advise deceleration. This visual aid allows drivers to navigate effectively without prior expertise in sim racing.
@@ VI-A Baselines
For LRSF, we follow the formulation in (7). This filter only intervenes when the system reaches the boundary of the safe set and does not consider the human operator’s input during interventions. We anticipate that such an abrupt approach which disregards human intent may confuse drivers and undermine comfort, thus increasing the risk of automation surprise. In contrast, our HCSF solves the optimization problem (11) to minimally modify the human action while still satisfying the Q-CBF constraint, which we hypothesize will enhance human agency, comfort, and overall satisfaction compared to LRSF. To ensure a fair comparison between our HCSF and LRSF, we employ the same neural approximation of the state–action safety value function ${\mathcal{Q}}(\cdot,\cdot)$ for both value-based safety filters. This ensures that both filters rely on the same safety monitor, while their interventions may differ (7), (11). In addition, each filter’s visual cues follow the same proportionality rule, maintaining consistency in how interventions are conveyed to human drivers.